# Expressive Power of Graph Neural Networks for (Mixed-Integer) Quadratic Programs

### v1

@@ Abstract.
Quadratic programming (QP) is the most widely applied category of problems in nonlinear programming. Many applications require real-time/fast solutions, though not necessarily with high precision. Existing methods either involve matrix decomposition or use the preconditioned conjugate gradient method. For relatively large instances, these methods cannot achieve the real-time requirement unless there is an effective preconditioner.
Recently, graph neural networks (GNNs) opened new possibilities for QP. Some promising empirical studies of applying GNNs for QP tasks show that GNNs can capture key characteristics of an optimization instance and provide adaptive guidance accordingly to crucial configurations during the solving process, or directly provide an approximate solution. Despite notable empirical observations, theoretical foundations are still lacking. In this work, we investigate the expressive or representative power of GNNs, a crucial aspect of neural network theory, specifically in the context of QP tasks, with both continuous and mixed-integer settings. We prove the existence of message-passing GNNs that can reliably represent key properties of quadratic programs, including feasibility, optimal objective value, and optimal solution. Our theory is validated by numerical results.
@@ 1. Introduction
Quadratic programming (QP) is an important type of optimization problem, with extensive applications across domains such as graph matching, portfolio optimization, and dynamic control [vogelstein2015fast, markowitz52, rockafellar1987linear]. The goal of QP is to minimize a quadratic objective function while satisfying specified constraints. These constraints can vary, leading to different subcategories of QP. When all the constraints are linear, we call a QP problem a linearly constrained quadratic program (LCQP). When they also involve quadratic inequalities, we call the problem a quadratically constrained quadratic program (QCQP). Furthermore, if the problem requires some variables to be integers, we call it mixed-integer QP. In this study, we focus on LCQP and its mixed-integer variant MI-LCQP.
In many real-world applications, finding solutions quickly is crucial, even if they are not perfectly precise. For example, in transportation systems, such as ride-hailing platforms like Uber or Lyft, matching drivers with passengers requires quick decision-making to minimize waiting times, even if the optimal solution is not attained. Similarly, in financial trading, algorithms must swiftly adjust investment portfolios in response to market changes, even if it is not the most optimal move.
@@
Unfortunately, existing methods for solving QP often rely on some computationally expensive techniques such as matrix decomposition and the preconditioned conjugate gradient method. For instance, matrix decomposition techniques like LU decomposition typically require $\mathcal{O}(n^{3})$ operations for a matrix with size $n\times n$ [golub2013matrix], although more advanced algorithms can achieve lower complexities. Similarly, the preconditioned conjugate gradient method involves $\mathcal{O}(n^{2})$ operations per iteration, and a high condition number of the matrix can lead to slow convergence or numerical instability [shewchuk1994introduction]. These considerations underscore the clear need for novel techniques to address the demands of real-time applications.
@@
Machine learning (ML) brings new chances to QP. Recent research indicates that deep neural networks (DNNs) can significantly improve the efficiency of the QP solving process. Based on the role of DNNs in the solving process, these studies can be broadly categorized into two classes:
@@
  • Type I: DNNs are used to accelerate an existing QP solver by processing QP instances along with relevant intermediate solving information. The DNNs then generate adaptive configurations tailored to the specific instance and context, optimizing the QP solving process. Success in this application relies on the DNN’s ability to capture in-depth features of QP instances and provide customized guidance to the solver [bonami2018learning, bonami2022classifier, ichnowski2021accelerating, getzelman2021learning, jung2022learning, king2024metric].
@@
  • Type II: DNNs replace or warm-start a QP solver. Here, DNNs take in a QP instance and directly output an approximate solution. This approximate solution can be used directly or as an initial solution for further refinement by a QP solver [nowak2017note, chen2018approximating, karg2020efficient, wang2020combinatorial, wang2020learning, wang2021neural, qu2021adaptive, gao2021deep, bertsimas2022online, liu2022revocable, sambharya2023end, pei2023reinforcement, tan2024ensemble].
Among the various types of DNNs, this paper focuses on graph neural networks (GNNs) due to their properties that align well with QP tasks. GNNs [scarselli2008graph] are defined on graphs and have been applied widely in many areas such as recommender systems, traffic management, and chemistry [wu2020comprehensive, zhou2020graph]. By conceptualizing QP problems as graphs, where all pertinent information including coefficients and boundaries are encoded into the graph’s attributes, GNNs can efficiently process and solve these QP tasks [nowak2017note, wang2020learning, wang2021neural, qu2021adaptive, gao2021deep, tan2024ensemble, jung2022learning]. An example of such a graph representation is illustrated in Figure 1. This approach leverages key advantages of GNNs: they adapt naturally to different sizes of graphs, allowing the same model to be applied to various QP problems without modification, and they are permutation invariant, ensuring that the output remains consistent regardless of the order of nodes. These features make GNNs particularly suitable for QP tasks, including their simplified forms like linear programming (LP).
$\min~{}\frac{1}{2}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}&{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}&{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}\begin{bmatrix}1&0.1&\\ 0.1&1&\\ &&1\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}+\begin{bmatrix}1&2&3\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}$s.t.
$2{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}$ $+$ ${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $\leq 5$
${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $+$ $3{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}$ $\geq 0$
$0\leq{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}},{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}},{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\leq 1$
$w_{1}$$(1,0,1)$$w_{2}$$(2,0,1)$$w_{3}$$(3,0,1)$$v_{1}$$(5,\leq)$$v_{2}$$(0,\geq)$21130.10.1111
Figure 1. An illustrative example of LCQP and its graph representation.
However, despite notable empirical results, a systematic understanding of GNN for QP is still lacking. To thoroughly understand its pros and cons, some critical questions must be addressed:
  • (Existence). Are there GNNs that can either capture the essential characteristics of QPs or provide approximate solutions? This question is named the expressive power of GNNs.
  • (Trainability). If such GNNs exist, can we find them? The process of finding such GNNs is named training, which involves gathering data, creating a method to measure success or failure (a loss function), and then refining the GNN to reduce the loss function.
  • (Generalization). Can a trained GNN perform effectively on QP instances it has not previously encountered? This concerns the generalization ability of GNNs.
@@
This paper primarily addresses the first question about the expressive power. For Type I applications, we investigate whether GNNs can accurately map a QP to its crucial features, focusing on feasibility and the optimal objective value. For Type II, we examine whether GNNs can map a QP to one of its optimal solutions. Formally, the question motivating this paper is:
(1.1) Are there GNNs that can accurately predict the feasibility,
optimal objective value, and an optimal solution of a QP?
The literature has explored the expressive capabilities of GNNs on general graph tasks [xu2019powerful, azizian2020expressive, geerts2022expressiveness, zhang2023expressive, li2022expressive, sato2020survey] and their ability to approximate continuous functions on graphs [azizian2020expressive, geerts2022expressiveness]. However, significant gaps remain in fully understanding how these capabilities relate to QP, as the connections between QP features (such as feasibility and optimal objective value) and graph properties have not been thoroughly established.
Contributions of this paper include:
  • (GNN for LCQP). We provide an affirmative answer to question (1.1), establishing a theoretical foundation for using GNNs for LCQP, across both Type I and II applications.
  • (GNN for MI-LCQP). In the case of MI-LCQP, our findings generally suggest a negative answer to question (1.1). However, we identify specific, precisely defined subclasses of MI-LCQP where GNNs can accurately predict feasibility, boundedness, and an optimal solution.
  • (Experimental Validation). We conduct experiments that directly validate the above results.
@@ 2. Preliminaries
This section introduces foundational concepts and preliminary definitions. We focus on linearly constrained quadratic programming (LCQP), which is formulated as follows:
@@
where $Q\in\mathbb{R}^{n\times n}$, $c\in\mathbb{R}^{n}$, $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, $l\in(\mathbb{R}\cup\{-\infty\})^{n}$, $u\in(\mathbb{R}\cup\{+\infty\})^{n}$, and {,=,}$\circ\in\{\leq,=,\geq\}^{m}$. In this paper, we assume $Q$ is symmetric and positive semidefinite, then the problem (2.1) is a convex program.
@@
Basic concepts of LCQPs. An $x$ that satisfies all the constraints of (2.1) is named a feasible solution. The set of all feasible solutions, defined as $X=:\{x\in\mathbb{R}^{n}:Ax\circ b,\ l\leq x\leq u\}$, is referred to as the feasible set. The LCQP is considered feasible if this set is non-empty; otherwise, it is infeasible. The value of $\frac{1}{2}x^{\top}Qx+c^{\top}x$ is named the objective (function) value. Its infimum across the feasible set is termed the optimal objective value. If this infimum is unattainable ($\infty$- ∞), suggesting the objective value could indefinitely decrease, the LCQP is deemed unbounded. Conversely, when the optimal objective value is realized, the corresponding $x$ is identified as an optimal solution. According to [eaves1971quadratic], a feasible and bounded LCQP must yield an optimal solution, though it might not be unique.
Graph representation of LCQPs. We introduce a graph structure, termed the LCQP-graph $G_{\textup{LCQP}}=(V,W,A,Q,H_{V},H_{W})$, that encodes all the elements of a LCQP (2.1). Particularly,
  • The graph contains two distinct types of nodes. Nodes in $V=\{1,2,\dots,m\}$, labeled as $i$, represent the $i$-th constraint and are called constraint nodes. Nodes in $W=\{1,2,\dots,n\}$, labeled as $j$, represent the $j$-th variable and are known as variable nodes. The union set $V\cup W$ includes all the vertices of the entire LCQP-graph $G_{\textup{LCQP}}$.
  • The graph comprises two distinct edge types. An edge connects $i\in V$ to $j\in W$ if $A_{ij}$ is nonzero, with $A_{ij}$ serving as the edge weight. Similarly, the edge between nodes $j,j^{\prime}\in W$ exists if $Q_{jj^{\prime}}\neq 0$, with $Q_{jj^{\prime}}$ as the edge weight. Self loops ($j=j^{\prime}$) are permitted.
  • Attributes/features $v_{i}=(b_{i},\circ_{i})$ are attached to the $i$-th constraint node for $i\in V$. The collection of all such attributes is denoted as $H_{V}=(v_{1},v_{2},\dots,v_{m})$.
  • Attributes/features $w_{j}=(c_{j},\ell_{j},u_{j})$ are attached to the $j$-th variable node for $j\in W$. The collection of all such attributes is denoted as $H_{W}=(w_{1},w_{2},\dots,w_{n})$.
@@
Such a representation is illustrated by an example shown in Figure 1. To the best of our knowledge, this particular representation is only detailed in [jung2022learning], yet it forms the foundation or core module for numerous related studies. For instance, removing nodes in $V$ and their associated edges reduces the graph into the assignment graph used in graph matching problems [nowak2017note, wang2020learning, wang2021neural, qu2021adaptive, gao2021deep, tan2024ensemble]. In these cases, the linear constraints $Ax\circ b$ are typically bypassed by applying the Sinkhorn algorithm to ensure that $x$ meets these constraints. Another scenario involves LP and MILP: removing edges associated with $Q$ simplifies the graph to a bipartite structure, which reduces the LCQP to an LP [chen2022representing-lp, fan2023smart, liu2024learning, qian2024exploring]. Further, by incorporating an additional node feature—an approach detailed in Section 4—this bipartite graph is also capable of representing MILP [gasse2019exact, chen2022representing-milp, nair2020solving, gupta2020hybrid, shen2021learning, gupta2022lookback, khalil2022mip, paulus2022learning, scavuzzo2022learning, liu2022learning, huang2023searching]. Notably, the body of work on MILP using this approach extends well beyond the cited references, indicating its robust and growing interest. The demonstrated empirical success of utilizing graph representations and GNNs in various optimization tasks underpins the motivation for this paper: we aim to explore the theoretical foundations of these methodologies.
@@
GNNs for solving LCQPs. Building on the established concepts, we present message-passing graph neural networks (hereafter referred to simply as GNNs) tailored for LCQPs using LCQP-graphs. These GNNs take in an LCQP-graph $G_{\textup{LCQP}}$ (including all the node features and edge weights) as input and update node features sequentially across layers via a message-passing mechanism. Initially, node features $s_{i}^{0},t_{j}^{0}$ are computed using embedding mappings $f_{0}^{V},f_{0}^{W}$:
@@
The architecture includes $L$ standard message-passing layers where each layer (where $1\leq l\leq L$) updates node features by locally aggregating neighbor information:
  • $s_{i}^{l}=f_{l}^{V}\big{(}s_{i}^{l-1},\sum_{j\in W}A_{ij}g_{l}^{W}(t_{j}^{l-1})\big{)}$ for $i\in V$, and
  • $t_{j}^{l}=f_{l}^{W}\big{(}t_{j}^{l-1},\sum_{i\in V}A_{ij}g_{l}^{V}(s_{i}^{l-1}),\sum_{j^{\prime}\in W}Q_{jj^{\prime}}g_{l}^{Q}(t_{j^{\prime}}^{l-1})\big{)}$ for $j\in W$.
Finally, there are two types of output layers. For applications where the GNN maps LCQP-graphs to a singular real value, such as evaluating properties like feasibility of the LCQP, a graph-level output layer is employed that computes a single real number encompassing the entire graph:
  • $y=r_{1}\big{(}\sum_{i\in V}s_{i}^{L},\sum_{j\in W}t_{j}^{L}\big{)}\in\mathbb{R}$.
Alternatively, if the GNN is required to map the LCQP-graph to a vector $y\in\mathbb{R}^{n}$, assigning a real number to each variable node as its output (as is typical in applications where GNNs are used to predict solutions), then a node-level output should be utilized. This output layer computes the value for the $j$-th output as follows:
  • $y_{j}=r_{2}\big{(}\sum_{i\in V}s_{i}^{L},\sum_{j\in W}t_{j}^{L},t_{j}^{L}\big{)}$.
In our theoretical analysis, we assume all the mappings $f_{l}^{V},f_{l}^{W}\ (0\leq l\leq L)$, $g_{l}^{V},f_{l}^{W},g_{l}^{Q}\ (1\leq l\leq L)$, and $r_{1},r_{2}$ to be continuous. In practice, these continuous mappings are learned from data. We aim to find mappings that enable all the LCQP-graphs $G_{\textup{LCQP}}$ from a dataset to be mapped accurately to their desired outputs $y$. To achieve this, we parameterize these mappings using multilayer perceptrons (MLPs) and optimize them within the parametric space. Based on all the concepts above, we present some definitions.
@@
The set of all LCQP-graphs, denoted as $\mathcal{G}^{m,n}_{\textup{LCQP}}$, comprises graphs with $m$ constraints and $n$ variables, where the matrix $Q$ is symmetric and positive semidefinite.
@@
The collection of all message-passing GNNs, denoted as $\mathcal{F}_{\textup{LCQP}}$ for graph-level outputs (or $\mathcal{F}_{\textup{LCQP}}^{W}$ for node-level outputs), consists of networks constructed using continuous mappings $f_{l}^{V},f_{l}^{W}\ (0\leq l\leq L)$, $g_{l}^{V},f_{l}^{W},g_{l}^{Q}\ (1\leq l\leq L)$, and $r_{1}$ (or $r_{2}$).
Note that the input graph size for GNNs within $\mathcal{F}_{\textup{LCQP}}$ and $\mathcal{F}_{\textup{LCQP}}^{W}$ is unspecified, as the functions $f_{l}^{V},f_{l}^{W}\ (0\leq l\leq L)$, $g_{l}^{V},f_{l}^{W},g_{l}^{Q}\ (1\leq l\leq L)$, and $r_{1}$ (or $r_{2}$) are independent of $m,n$. This independence highlights a key advantage of GNNs discussed in Section 1: their adaptability to various graph sizes, allowing the same model to be consistently applied across different QPs.

### v2

@@ Abstract.
Quadratic programming (QP) is the most widely applied category of problems in nonlinear programming. Many applications require real-time/fast solutions, though not necessarily with high precision. Existing methods either involve matrix decomposition or use the preconditioned conjugate gradient method. For relatively large instances, these methods cannot achieve the real-time requirement unless there is an effective preconditioner. Recently, graph neural networks (GNNs) opened new possibilities for QP. Some promising empirical studies of applying GNNs for QP tasks show that GNNs can capture key characteristics of an optimization instance and provide adaptive guidance accordingly to crucial configurations during the solving process, or directly provide an approximate solution. However, the theoretical understanding of GNNs in this context remains limited. Specifically, it is unclear what GNNs can and cannot achieve for QP tasks in theory. This work addresses this gap in the context of linearly constrained QP tasks. In the continuous setting, we prove that message-passing GNNs can universally represent fundamental properties of convex quadratic programs, including feasibility, optimal objective values, and optimal solutions. In the more challenging mixed-integer setting, while GNNs are not universal approximators, we identify a subclass of QP problems that GNNs can reliably represent.
ZC, XC, and JL contribute equally. Corresponding author: Jialin Liu, jialin.liu@ucf.edu.
The work of ZC is supported in part by the National Science Foundation via grant DMS-2509011.
@@ 1. Introduction
Quadratic programming (QP) is an important type of optimization problem with applications [vogelstein2015fast, markowitz52, rockafellar1987linear]. It aims to minimize a quadratic objective function while satisfying specified constraints. When all the constraints are linear, we call a QP problem a linearly constrained quadratic program (LCQP). When they also involve quadratic inequalities, we call the problem a quadratically constrained quadratic program (QCQP). When some variables are restricted to integers, the problem becomes a mixed-integer QP. This study focuses on LCQP and its mixed-integer variant, MI-LCQP.
In many applications, finding solutions quickly is prioritized over perfect precision. For instance, ride-hailing platforms like Uber or Lyft require quick driver-passenger matching to reduce wait times, even without optimal solutions. Similarly, financial trading algorithms must rapidly adjust portfolios to market changes, prioritizing speed over optimality.
@@
Unfortunately, existing methods for QP often rely on computationally expensive techniques such as matrix decomposition and the preconditioned conjugate gradient method (PCG). For example, LU decomposition typically requires $\mathcal{O}(n^{3})$ operations for a $n\times n$ matrix [golub2013matrix], though advanced algorithms can achieve lower complexities. The PCG method requires $\mathcal{O}(n^{2})$ operations per iteration, with slow convergence for ill-conditioned matrices [shewchuk1994introduction]. These challenges highlight the need for novel approaches to meet real-time application demands.
@@
Machine learning brings new chances to QP. Recent research shows that deep neural networks (DNNs) can significantly improve the efficiency when solving QP. Based on DNNs’ role, these studies can be categorized as:
@@
(Type I). DNNs generate adaptive configurations for QP solvers, tailored to specific QP instances, thereby accelerating the solving process. [bonami2018learning, bonami2022classifier, ichnowski2021accelerating, getzelman2021learning, jung2022learning, king2024metric]. This approach requires DNNs to capture in-depth features of QP instances and provide customized guidance to the solver.
@@
(Type II). DNNs replace or warm-start a QP solver. Here, DNNs take in a QP and directly output an approximate solution. These solutions can be used as final outputs or as initial guesses to accelerate QP solvers [nowak2017note, chen2018approximating, karg2020efficient, wang2020combinatorial, wang2020learning, wang2021neural, qu2021adaptive, gao2021deep, bertsimas2022online, liu2022revocable, sambharya2023end, pei2023reinforcement, tan2024ensemble].
GNNs.
Among the various types of DNNs, this paper focuses on graph neural networks (GNNs)  [scarselli2008graph]. By conceptualizing QPs as graphs (Figure 1), GNNs can be applied and efficiently handle these tasks [nowak2017note, wang2020learning, wang2021neural, qu2021adaptive, gao2021deep, tan2024ensemble, jung2022learning]. For instance, [wang2021neural] use GNNs to solve Lawler’s QAP [lawler1963quadratic], while [wang2019learning, Yu2020Learning] apply GNNs to Koopman-Beckmann’s QAP [loiola2007survey]. They exploit key strengths of GNNs: adaptability to varying graph sizes, allowing the same model applied to various QPs, and permutation invariance, ensuring consistent outputs regardless of node order.
Expressive power
Despite their notable advantages, GNNs face fundamental limitations. As [xu2019powerful] pointed out, GNNs’ expressive power is limited: they are not universal approximators for all graph-based functions. Here, expressive power, a core concept in deep learning theory, measures the existence of neural networks under a given structure that can approximate (or represent) a broad class of functions, and universal approximation guarantees that a model can approximate any functions within its domain.
The contrast between the successful empirical applications of GNNs and their theoretical limitations reveals a significant gap. However, in practice, GNNs do not need to approximate all possible functions but only specific, meaningful mappings relevant to QPs. This leads to the key question motivating this work: Can GNNs, despite their limitations for general graph-based mappings, exhibit sufficient expressive power to predict the key properties of QPs?
@@
To address this question, we focus on two types of functions according to the nature of the application. For Type I applications, we investigate whether GNNs can accurately map a QP to its critical features, focusing on the feasibility and optimal objective value. For Type II, we explore whether GNNs can map a QP to one of its optimal solutions. Therefore, we ask:
(1.1) Can GNNs accurately predict the feasibility, optimal
objective value, and an optimal solution of a QP?
The literature has explored the expressive powers of GNNs on general graph tasks [xu2019powerful, azizian2020expressive, geerts2022expressiveness, zhang2023expressive, li2022expressive, sato2020survey]. However, significant gaps remain in understanding how these results relate to QP. The most relevant works [chen2022representing-lp, chen2022representing-milp] investigate the expressive power of GNNs for (mixed-integer) linear programs (MILPs), but their analysis highly depends on the linear structure and does not cover nonlinear programs like QP. A concurrent study by [wu2024representing] investigates GNN representations for quadratically constrained quadratic programs (QCQPs) using a tripartite graph structure. While both studies address convex LCQPs, the methodologies and focus differ: [wu2024representing] consider general QCQPs, including quadratic constraints, and use a tripartite graph representation, whereas our work focuses on (MI-)LCQPs with a bipartite graph construction. Our analysis also covers mixed-integer settings and proposes an alternative approach to handling quadratic constraints, yielding partially overlapping but distinct results.
Contributions
As several studies have empirically shown that incorporating GNNs can significantly enhance the performance of QP solvers, this paper aims to theoretically analyze the expressive power of GNNs in such tasks, identify potential areas for improvement, and highlight key considerations. Specifically, the contributions of this paper include:
  • (GNN for LCQP). In the continuous setting, where all variables are allowed to take fractional numbers, we provide an affirmative answer to question (1.1) with convexity and present a nonconvex counterexample.
  • (GNN for MI-LCQP). In mixed-integer settings, where some variables must be integers, we provide counterexamples showing that GNNs cannot universally solve all tasks, giving a negative answer to question (1.1).
  • Despite this limitation, we identify specific, precisely defined subclasses of MI-LCQP where GNNs succeed. Importantly, we also present criteria for determining whether an MI-LCQP belongs to this subclass, which can be efficiently verified numerically.
@@ 2. Preliminaries
We focus on linearly constrained quadratic programming (LCQP), which is formulated as follows:
@@
where $Q\in\mathbb{R}^{n\times n}$, $c\in\mathbb{R}^{n}$, $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, $l\in(\mathbb{R}\cup\{-\infty\})^{n}$, $u\in(\mathbb{R}\cup\{+\infty\})^{n}$, and $\circ\in\{\leq,=,\geq\}^{m}$. In this paper, we always assume that $Q$ is symmetric.
Basic concepts of LCQPs
@@
An $x$ satisfying all constraints of (2.1) is named a feasible solution. The set of all feasible solutions, $X=:\{x\in\mathbb{R}^{n}:Ax\circ b,\ l\leq x\leq u\}$, is referred to as the feasible set. The LCQP is feasible if this set is non-empty; otherwise, it is infeasible. The value of $\frac{1}{2}x^{\top}Qx+c^{\top}x$ is the objective value. Its infimum across $X$ is termed the optimal objective value. If this infimum is -$\infty$ (the objective value could indefinitely decrease), the LCQP is named unbounded. A feasible and bounded LCQP must yield an optimal solution [eaves1971quadratic].
$w_{1}=$$(c_{1},l_{1},+\infty)$$w_{2}=$$(c_{2},l_{2},+\infty)$$w_{3}=$$(c_{3},-\infty,u_{3})$$v_{1}=$$(b_{1},\leq)$$v_{2}=$$(b_{2},\geq)$$A_{11}$$A_{12}$$A_{22}$$A_{23}$$Q_{12}$$Q_{21}$$Q_{11}$$Q_{22}$$Q_{33}$
$\min$ $c_{1}~{x_{1}}$ ++ $c_{2}~{x_{2}}$ ++ $c_{3}~{x_{3}}$ $\frac{1}{2}[{x_{1}}~{x_{2}}~{x_{3}}]\par\begin{bmatrix}Q_{11}&Q_{12}&\\ Q_{21}&Q_{22}&\\ &&Q_{33}\end{bmatrix}\par\begin{bmatrix}{x_{1}}\\ {x_{2}}\\ {x_{3}}\end{bmatrix}$
s.t. $x_{1}$} $\geq l_{1},$ $x_{2}$} $\geq l_{2},$ $x_{3}$} $\leq u_{3}$
$A_{11}~{x_{1}}$ ++ $A_{12}~{x_{2}}$ $\leq b_{1}$
$A_{22}~{x_{2}}$ ++ $A_{23}~{x_{3}}$ $\geq b_{2}$
Figure 1. LCQP and its graph representation: All the information from the LCQP is fully encoded within the node or edge attributes.
Graph representation
We present a graph structure, termed the LCQP-graph, $G_{\textup{LCQP}}=(V,W,A,Q,H_{V},H_{W})$, that encodes all the elements of a LCQP. The graph contains two distinct types of nodes: Nodes in $V=\{1,2,\dots,m\}$, labeled as $i$, represent the $i$-th constraint and are called constraint nodes; Nodes in $W=\{1,2,\dots,n\}$, labeled as $j$, represent the $j$-th variable and are known as variable nodes. An edge connects $i\in V$ to $j\in W$ if $A_{ij}$ is nonzero, with $A_{ij}$ serving as the edge weight. Similarly, the edge between nodes $j,j^{\prime}\in W$ exists if $Q_{jj^{\prime}}\neq 0$, with $Q_{jj^{\prime}}$ as the edge weight. Self loops ($j=j^{\prime}$) are permitted. Attributes (or features) $v_{i}=(b_{i},\circ_{i})$ are attached to the $i$-th constraint node for $i\in V$. The collection of all such attributes is denoted as $H_{V}=(v_{1},v_{2},\dots,v_{m})$. Attributes $w_{j}=(c_{j},\ell_{j},u_{j})$ are attached to the $j$-th variable node for $j\in W$ and their collection is denoted as $H_{W}=(w_{1},w_{2},\dots,w_{n})$.
@@
Such a representation, illustrated by Figure 1, is a fundamental and “minimal” approach in the sense that every entry in $(A,b,c,Q,l,u,\circ)$ is used exactly once. While this particular representation is only detailed in [jung2022learning], it forms the foundation of numerous related studies. For instance, removing nodes in $V$ and their associated edges reduces the graph into the assignment graph used in graph matching problems [nowak2017note, wang2020learning, wang2021neural, qu2021adaptive, gao2021deep, tan2024ensemble]. Removing edges associated with $Q$ simplifies the graph to a bipartite structure and reduces LCQP to LP [chen2022representing-lp, fan2023smart, liu2024learning, qian2024exploring]. By adding an extra node feature, an approach detailed in Section 4, this graph can also express mixed-integer programs [gasse2019exact, chen2022representing-milp, nair2020solving, gupta2020hybrid, shen2021learning, gupta2022lookback, khalil2022mip, paulus2022learning, scavuzzo2022learning, liu2022learning, huang2023searching, wang2024digmilp].
GNNs for LCQPs
@@
Given the graph representation, we present message-passing graph neural networks (hereafter referred to simply as GNNs) for LCQPs. They take in an LCQP-graph $G_{\textup{LCQP}}$ (including the node and edge attributes) and update node attributes sequentially across layers via a message-passing mechanism. Initially, node attributes are updated separately using embedding mappings $f_{0}^{V},f_{0}^{W}$:
@@
The architecture includes $L$ standard message-passing layers where each layer (where $1\leq l\leq L$) updates node attributes by locally aggregating neighbor information:
$s_{i}^{l}=$ $f_{l}^{V}\left(s_{i}^{l-1},\sum_{j\in\mathcal{N}_{i}^{W}}g_{l}^{W}(t_{j}^{l-1},A_{ij})\right),$
$t_{j}^{l}=$ $f_{l}^{W}\left(t_{j}^{l-1},\sum_{i\in\mathcal{N}_{j}^{V}}g_{l}^{V}(s_{i}^{l-1},A_{ij}),\sum_{j^{\prime}\in\mathcal{N}_{j}^{W}}g_{l}^{Q}(t_{j^{\prime}}^{l-1},Q_{jj^{\prime}})\right),$
where $f_{l}^{V},f_{l}^{W},g_{l}^{V},g_{l}^{W},g_{l}^{Q}$ are trainable local updates in GNNs and $\mathcal{N}_{i}^{W}=\{j\in W:A_{ij}\neq 0\}$, $\mathcal{N}_{j}^{V}=\{j\in V:A_{ij}\neq 0\}$, and $\mathcal{N}_{j}^{W}=\{j^{\prime}\in W:Q_{jj^{\prime}}\neq 0\}$ are the sets of neighbors. Finally, there are two types of output layers. For applications where the GNN maps LCQP-graphs to a real value, such as evaluating properties like feasibility of LCQP, a graph-level output layer is employed that computes a single real number encompassing the entire graph:
$y=r_{1}\left(\sum_{i\in V}s_{i}^{L},\sum_{j\in W}t_{j}^{L}\right)\in\mathbb{R}.$
Alternatively, if the GNN is required to map the LCQP-graph to a vector $y\in\mathbb{R}^{n}$, assigning a real number to each variable node as its output (as is typical in applications where GNNs are used to predict solutions), then a node-level output should be utilized:
$y_{j}=r_{2}\left(\sum_{i\in V}s_{i}^{L},\sum_{j\in W}t_{j}^{L},t_{j}^{L}\right),\quad\textup{for $j\in W$.}$
Here, $r_{1}$ and $r_{2}$ are trainable output functions. In our theoretical analysis, we assume all the mappings $f_{l}^{V},f_{l}^{W}\ (0\leq l\leq L)$, $g_{l}^{V},f_{l}^{W},g_{l}^{Q}\ (1\leq l\leq L)$, and $r_{1},r_{2}$ to be continuous. In practice, these continuous mappings are usually parameterized by multilayer perceptrons (MLPs) and their parameters are learned from data.
@@
The set of all LCQP-graphs, denoted as $\mathcal{G}^{m,n}_{\textup{LCQP}}$ [^1] The space $\mathcal{G}^{m,n}_{\textup{LCQP}}$ is equipped with the subspace topology induced from the product space $\big\{(A,b,c,Q,l,u,\circ):A\in\mathbb{R}^{m\times n},b\in\mathbb{R}^{m},c\in\mathbb{R}^{n},Q\in\mathbb{R}^{n\times n},l\in(\mathbb{R}\cup\{-\infty\})^{n},u\in(\mathbb{R}\cup\{+\infty\})^{n},\circ\in\{\leq,=,\geq\}^{m}\big\}$. All Euclidean spaces have standard Eudlidean topologies, discrete spaces are equipped with the discrete topology, and their unions are disjoint unions., comprises graphs with $m$ constraints and $n$ variables, where $Q$ is symmetric.
@@
The collection of all GNNs, denoted as $\mathcal{F}_{\textup{LCQP}}$ for graph-level outputs (or $\mathcal{F}_{\textup{LCQP}}^{W}$ for node-level outputs), consists of all GNNs constructed using continuous mappings $f_{l}^{V},f_{l}^{W}\ (0\leq l\leq L)$, $g_{l}^{V},f_{l}^{W},g_{l}^{Q}\ (1\leq l\leq L)$, and $r_{1}$ (or $r_{2}$).